OpenStudy (anonymous):

Divide and simplify. w^2-1/9w+9\[\div\]w-1/3

jhonyy9 (jhonyy9):

(w-1)(w+1)/(9(w+1)) = (w-1)/9

jhonyy9 (jhonyy9):

ah there is again divide

jhonyy9 (jhonyy9):

so

jhonyy9 (jhonyy9):

(w-1)/9 * 3/(w-1) = 1/3

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

when you need divide two fractions you need multiply first farction by invers of secondly

jhonyy9 (jhonyy9):

than you get (w-1)/9 * 3/(w-1) (w-1) 3 ----- * ----- = 9 (w-1)